OpenStudy (anonymous):

find the equation of the line that passes through the point (-3,3), and that is perpendicular to the line with the equation y=1/3x+5/3

OpenStudy (anonymous):

Since I can't just post the answer, I'll walk you through the problem okay?

OpenStudy (anonymous):

preciate that

OpenStudy (anonymous):

you know what a perpendicular line is right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Its an intersecting line. + basically the opposite of parallel lines =

OpenStudy (anonymous):

ok \

OpenStudy (anonymous):

Perpendicular lines have a line with positive slope (so it's an increasing line), then the perpendicular line must have negative slope (because it will be a decreasing line). So perpendicular slopes have opposite signs (+) (-) . The other "opposite" thing with perpendicular slopes is that their values are reciprocals

OpenStudy (anonymous):

y=-3x?

OpenStudy (anonymous):

exactly then when you graph make sure it passes through the given points

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

You could honestly start at 3,3 and use the slope to -3/1 to graph it. After it is just a matter of plugging in the numbers for the variables y= -3/1x+3

OpenStudy (anonymous):

i dont need to graph it it just wants me to give the equation

OpenStudy (anonymous):

the graph is only if you want to check your answer

OpenStudy (anonymous):

so whats the equation

OpenStudy (anonymous):

whats the anwer

OpenStudy (anonymous):

y= -3/1x+3

OpenStudy (anonymous):

shoudnt it just be y=-3x+3

OpenStudy (anonymous):

yes sorry typing too fast

OpenStudy (anonymous):

??

OpenStudy (anonymous):

your right its y= -3/1x+3

OpenStudy (anonymous):

im saying shondt it just be y=-3x+3 instead of y=-3/1x+3

OpenStudy (anonymous):

its the same equation - 3 = - 3/1 doesn't matter if its on fraction form or not